OpenStudy (anonymous):

Intergration please solve intergral tanx^3 dx

OpenStudy (amistre64):

odd degree trig functions; I think its easier to split them up into groups of something...

OpenStudy (anonymous):

Re-write it as ∫tan^2(x)*tan(x) dx ---> ∫(sec^2(x)-1)*tan(x) dx = ∫sec^2(x)tan(x) - tan(x) dx. Let u=tan(x).

OpenStudy (anonymous):

well, it is quite simple actually, just wait a minute and sstarica will show you

OpenStudy (anonymous):

re-write it and you'll get :\[\int\limits_{}^{}sinx^3/cosx^3 dx\] use u substitution for cosx^3 ^_^ give it a try :)

OpenStudy (anonymous):

LOL andy :D

OpenStudy (amistre64):

Here, use this :)

OpenStudy (anonymous):

oh, and ssarica, gratz on making a Superstar, I guess I cant call you starlet anymore

OpenStudy (anonymous):

lol, call me whatever you want and thank you ^_^, some of the fanning came from you :)

OpenStudy (anonymous):

yup, whole 4 fannings

OpenStudy (anonymous):

amistre! are you encouraging students on using calculators instead of thinking about it?!

OpenStudy (amistre64):

at 101 fans you get to be a "hero"....its a shame really, all that work just to become a sandwich....

OpenStudy (anonymous):

different times, different accounts, but same me

OpenStudy (anonymous):

:) I don't really care much about the fanning process

OpenStudy (amistre64):

the mr proffesor is more of an "exercise" in math than it is a calculator :)

OpenStudy (anonymous):

lol, anyway did you understand it deidre? ^_^

OpenStudy (anonymous):

@sstarica: I'm not sure if your method works, it ends up giving you -1/3 * ∫1/cos^5(x) dx. :/

OpenStudy (amistre64):

2x6=? 12 (green light)

OpenStudy (anonymous):

for reals? let me try

OpenStudy (anonymous):

Whoops, not quite..

OpenStudy (anonymous):

no lol , it works!

OpenStudy (anonymous):

Oh, sorry. Thought you typed something else ;P

OpenStudy (anonymous):

wait, you're right it doesn't work ^^" sorry for the mess again diedre, ignore my answer. Thank you Quantum ^_^

OpenStudy (anonymous):

deidre is away, geting some hotdogs for me and me

OpenStudy (anonymous):

Now let me check if my answer works...xD

OpenStudy (anonymous):

so dont worry, you have some time to clear the mess before he comes back

OpenStudy (anonymous):

\[du = -3x^2sinx^3dx\] which is wrong. you can't substitute it there, my dearest apologies ^^"

OpenStudy (anonymous):

Ah, let me expand my answer to make it clearer: ∫tan(x)sec^2(x) dx - ∫tan(x) dx. In the first integral, let u=tan(x), and in the second it integrates cleanly. :P

OpenStudy (anonymous):

why did you put tanxsec^2x? where did you get it from

OpenStudy (anonymous):

x^3 acts as an angle here

OpenStudy (anonymous):

and not the power of tan :)

OpenStudy (anonymous):

ok, guys, I am about to go to sleep, just going to workout abit and maybe eat something and i am off.

OpenStudy (anonymous):

same, I'm tired

OpenStudy (anonymous):

oh, nice, so we are going to beds together again

OpenStudy (anonymous):

anyway, solve the question and I'm off to bed, sorry ^_^

OpenStudy (anonymous):

I'm going to bed ALONE >_>

OpenStudy (anonymous):

I sayed BEDS with an S, not BED :D LOL, I knew it you have interesting thoughts in your head :)

OpenStudy (anonymous):

I got it from the relation tan^2(x) = sec^2(x) - 1. ∫tan^3(x)dx = ∫tan^2(x)*tan(x) dx = ∫(sec^2(x)-1)*tan(x) dx = ∫tan(x)sec^2(x) - tan(x) dx = ∫tan(x)sec^2(x) dx - ∫tan(x) dx. :P

OpenStudy (anonymous):

nope :)

OpenStudy (anonymous):

I never think of such things angoo =P

OpenStudy (anonymous):

yes but quantum the question is :\[\int\limits_{}^{}tanx^3\] and not : \[\int\limits_{}^{}\tan^3x\]

OpenStudy (anonymous):

hmm... well, I don't too, becouse things like those should be practice once married

OpenStudy (anonymous):

I agree :)

OpenStudy (anonymous):

quantum you copied the question wrong lol

OpenStudy (anonymous):

good, smart girl, not the "new era" type

OpenStudy (anonymous):

nah, I'm old fashioned ^_^ more classy

OpenStudy (anonymous):

wonderful, smart, classy, what else could one want....

OpenStudy (anonymous):

want what? ._.

OpenStudy (anonymous):

Erm...alright, then we'll see when deidre clears it up. :P

OpenStudy (anonymous):

like in, what else would I look for in a friend

OpenStudy (anonymous):

lol , I agree Quantum

OpenStudy (anonymous):

aww, angoo, you're soo sweet ^_^

OpenStudy (amistre64):

im sticking with my safety net....(S) x^2 dx..... :)

OpenStudy (anonymous):

not sweet, just relaxed and not trying to be someone I am not

OpenStudy (anonymous):

lol amistre! and yes you are sweet andy ^_^

OpenStudy (anonymous):

well, thanks, going to the dream world now, I think I will dream about beautiful land with happy little rabbits and dears and bears living together, and I will be ther too, we will play and have fun....

OpenStudy (anonymous):

lala land, lol, I'm off to bed too, good night ^_^

OpenStudy (anonymous):

and you are welcome :)

OpenStudy (anonymous):

night night :)

OpenStudy (anonymous):

sleep tight ~

OpenStudy (anonymous):

:) I was about to say "don't let the bugs bite" but it is getting old, so..

OpenStudy (anonymous):

lol, bye

OpenStudy (anonymous):

bye